# REvolve: Reward Evolution with Large Language Models for Autonomous Driving

### v1

@@ Abstract
Designing effective reward functions is crucial to training reinforcement learning (RL) algorithms. However, this design is non-trivial, even for domain experts, due to the subjective nature of certain tasks that are hard to quantify explicitly. In recent works, large language models (LLMs) have been used for reward generation from natural language task descriptions, leveraging their extensive instruction tuning and commonsense understanding of human behavior. In this work, we hypothesize that LLMs, guided by human feedback, can be used to formulate human-aligned reward functions. Specifically, we study this in the challenging setting of autonomous driving (AD), wherein notions of "good" driving are tacit and hard to quantify. To this end, we introduce REvolve, an evolutionary framework that uses LLMs for reward design in AD. REvolve creates and refines reward functions by utilizing human feedback to guide the evolution process, effectively translating implicit human knowledge into explicit reward functions for training (deep) RL agents. We demonstrate that agents trained on REvolve-designed rewards align closely with human driving standards, thereby outperforming other state-of-the-art baselines.
@@ 1 Introduction
Recently, large language models (LLMs) have shown impressive abilities in instruction following [29, 38], search and black-box optimization[28, 62, 65], self-reflection [32, 49], and code generation [6, 42]. Owing to their vast training and world knowledge [17, 19, 59], they can readily incorporate commonsense priors of human behavior. Together, these abilities make LLMs an ideal candidate for reward generation from natural language (NL) task descriptions and refining them using language-based feedback loops. Indeed, recent works like Language to Rewards [63] and Text2Reward [61] have utilized LLMs for this purpose.
@@
In particular, Eureka [30] utilizes LLMs to design dense and interpretable reward functions, achieving improved performance in dexterous manipulation tasks. It does so by generating multiple reward functions and greedily iterating with the best candidate over multiple iterations. We identify this greediness as a major limitation as it risks premature convergence and leads to a waste of resources (by preemptively discarding weaker reward functions). Moreover, Eureka relies on having access to fitness functions in order to assess the quality of learned behaviors. This is particularly challenging for complex tasks, such as AD, which require human evaluations.
@@
To this end, we introduce Reward Evolve (REvolve), a novel evolutionary framework using LLMs, specifically GPT-4, to output reward functions (as executable Python codes) for AD and evolve them based on human feedback. In particular, REvolve offers several key advantages:
Figure 1: An overview of REvolve. Given the task of AD and abstracted environment variables, a reward designer $G$ (LLM) outputs a population of reward functions $R\in\mathcal{R}$, each used to train an AD policy $\pi(R)$ in driving simulation. Then, we collect human preferences and natural language (NL) feedback on pairs of policy rollouts $\theta\sim\Theta_{\pi(R)}$ through a human user feedback interface. Policy (and thus, corresponding reward function) fitness $\sigma$ is calculated, and the fittest individuals, along with their NL feedback $\lambda$, are refined by $G$. The process leverages genetic programming for evolution. The flames symbolize trainable parameters.
(1) Framing reward design as a search problem. Based on evolutionary algorithms [2], REvolve evolves a population of reward functions by applying genetic operations such as mutation, crossover, and migration. This enhances genetic diversity and prevents premature convergence in complex search space (in comparison to greedy search), all without incurring additional computational costs compared to Eureka.
@@
(2) Utilizing human feedback to guide the search. Human preference data is directly mapped into fitness scores, effectively allowing humans to serve as fitness functions. Additionally, the framework incorporates qualitative natural language feedback, offering GPT-4 high-level insights into which aspects to prioritize while refining the reward functions. Together, this ensures that the reward functions – and thus the learned policies – align with human driving standards. Unlike inverse reinforcement learning (IRL) approaches, which require extensive demonstrations [46] or interventions [45] to infer human behavior indirectly, REvolve uses minimal feedback to directly refine reward functions.
@@
(3) Eliminating the need for additional reward model training. Unlike reinforcement learning with human feedback (RLHF) [9], which relies on a black-box reward model (trained on human preferences between outputs) to align LLMs with goals like safety and reliability [36], REvolve requires no extra models. REvolve-designed reward functions are interpretable (as executable Python code) compared to RLHF rewards, making it ideal for safety-critical applications, such as AD [8].
@@
Figure 1, provides an overview of REvolve for AD within a simulated environment.
@@ 2 Problem Formalization
Let us first formally introduce the reward design problem (RDP) [51], given by the tuple $\langle M,\mathcal{R},\pi,F\rangle$, where $M=(S,A,T)$ is the world model consisting of state space $S$, action space $A$, and transition function $T$. $\mathcal{R}$ is the set of all reward functions where each $R\in\mathcal{R}$ is a reward function that maps a state-action pair to a real-valued scalar $R:S\times A\rightarrow\mathbb{R}$. The tuple $(M,R)$ constitutes a Markov decision process [53]. Furthermore, $\pi:S\rightarrow\Delta A$ is a policy that maps a state to a distribution over actions, which is learned using a reward function $R$.
Moreover, the RDP includes a function $F$ that evaluates the fitness (i.e. quality) of the trained policies. Traditionally, $F$ maps policy rollouts, i.e. rollouts of agent-environment interactions, to real-valued scalars. Formally, we denote this by $F:\Omega(\Theta_{\pi})\rightarrow\mathbb{R}$, where $\Theta_{\pi}$ is the random variable modeling the probability distribution of rollouts induced by $\pi$ and $\Omega(\Theta_{\pi})$ is its sample space.
@@
In order to instantiate the RDP tuple for REvolve, we assume that we are given a reward designer $G:\mathcal{L}\rightarrow\mathcal{R}$ that maps from the set of all strings to the set of all reward functions $\mathcal{R}$. The objective now is to find a reward function $R\in\mathcal{R}$ that maximizes the expected fitness score of the policy $\pi$ trained on $R$, which we denote by $\pi(R)$. Formally we express the optimization problem as:
@@
An implicit assumption for reward design is the knowledge of observation and action variables in the environment. For real-world problems like AD, the assumption is supported by access to sensor data (such as speed from speedometers and distance covered from odometers), or by employing advanced vision-based models [7, 22, 64].
@@ 3 Reward Evolution
As the fitness function $F$ is based on implicit human understanding, it cannot be explicitly formulated and is therefore not amenable to gradient-based optimization. Therefore, we solve the RDP using a meta-heuristic approach through genetic programming (GP) [26]. Concretely, REvolve uses GP to output and iteratively refine a population of reward functions (individuals) by applying genetic operations such as mutation, crossover, and migration. In each iteration, we obtain new individuals, which are added to the population based on their fitness. This fitness is assessed through human evaluations of the policies trained with these reward functions – thus, guiding the evolutionary search.
Generally, GPs require significant domain expertise and rely on random initialization and perturbations for exploration, leading to scalability issues [37]. In contrast, LLMs can suggest changes that are intelligent and better aligned with the feedback [13]. Therefore, we combine the evolutionary search of REvolve with LLMs by using them as reward designers $G$ in the RDP.
We follow the island model of evolution [4] where we maintain multiple ($I$) sub-populations (or islands) in a reward database $D$. Island populations evolve by locally reproducing (mutation and crossover). Periodically, individuals migrate from one island to another. The island model helps enhance genetic diversity across the population and prevent premature convergence in complex search spaces [4]. Following existing works in reward design with LLMs [30, 61], we represent the reward functions as Python programs that seamlessly integrate with the GP framework. In what follows, we describe the four stages (initialization, reproduction, selection, termination) of our evolutionary framework. We give the formal description in Algorithm 1 – for brevity, we use $\pi$ instead of $\pi(R)$.
@@
10:         is_mutate $\sim\texttt{Bernoulli}(p_{m})$triangleright Sample operation
@@
25:             $D[P]\leftarrow\texttt{Update}(R,\pi,\sigma,\lambda)$triangleright Add reward to database subpopulation
@@ 3.1 Initialization
We start by creating a population of $K$ reward function individuals using GPT-4. This leverages GPT-4’s world knowledge [17] and its zero-shot instruction following abilities [38]. For each individual, we train a policy $\pi$, compute a fitness score $\sigma$, and obtain natural language feedback $\lambda\in\mathcal{L}$. The fitness score calculation and feedback collection are detailed in Section 3.3. Each individual and their respective policy, fitness score, and natural language feedback $(R,\pi,\sigma,\lambda)$ is then randomly assigned to a sub-population and stored in the reward database $D$.
@@ 3.2 Reproduction
Each successive generation of $K$ individuals is created by applying genetic operators (crossover and mutation), on the existing reward database individuals. We use GPT-4 as mutation and crossover operators with appropriate prompts. Due to space constraints, we provide prompt examples for mutation and crossover in Appendix D.
We begin by randomly selecting an operation – mutation with probability $p_{m}$ or crossover with probability $p_{c}=1-p_{m}$. Next, we sample a sub-population $P$ via weighted selection based on the average fitness scores $\sigma^{P}$ of each of the $I$ sub-populations. Subsequently, we sample parents (one for mutation, two for crossover) from $D[P]$ using weighted sampling based on individual fitness scores. These sampled individuals along with their natural language feedback $\lambda$ are used to prompt GPT-4 to create new individuals. These prompts direct GPT-4 to either modify a single reward component of the parent ($\texttt{Mutate}(G,D[P])$) or combine the best-performing reward components from two parents ($\texttt{Crossover}(G,D[P])$). Periodically, we enable the migration of individuals between sub-populations to enhance genetic diversity. Mutation and crossover operations are illustrated in Figure 2.
Figure 2: Illustration of how GPT-4 applies mutation and crossover to reward functions. Mutation (left): shows the modification of the "smoothness reward" component. A red ‘-’ sign indicates the line removed from the parent reward function, while a green ‘+’ sign indicates the line added to the new reward function. Crossover (right): demonstrates how parent reward functions are combined to create a child reward function, incorporating the most effective reward components from each parent.
@@ 3.3 Selection
Newly reproduced individuals (from the previous stage) are retained based on fitness scores, following a survival of the fittest approach. A fitness function $F$, tailored to the specific problem, evaluates the quality of each individual. The process involves (1) training a policy $\pi(R)$ until convergence, (2) sampling rollouts $\theta\sim\Theta_{\pi(R)}$, and (3) using the fitness function to obtain a tuple of (real-valued) fitness score and (string) NL feedback $(\sigma,\lambda)\coloneqq F(\theta)$. Individuals are retained if they contribute to an increase in the average fitness of their respective sub-populations $\sigma\geq\sigma^{P}$ where $P$ denotes the sub-population that the individual is added to [^1] To ensure genetic diversity, we add individuals based on surpassing the average and not maximum scores.. This ensures that the average fitness of the sub-population is monotonically increasing. In what follows, we detail the process of obtaining the tuple $(\sigma,\lambda)$ from the fitness function.
To compute the fitness score $\sigma$, we ask human evaluators to judge policy rollouts $\theta$ from different policies in a pairwise fashion. As shown by the "Human Feedback Interface" in Figure 1, these rollouts are presented as video segments $\approx$ 30-40 seconds long. The evaluators choose their preferred segment (ties are also possible). We then apply an Elo rating system [10], which transforms pairwise assessments into fitness scores $\sigma$. This avoids the need to learn a separate reward model, cf. RLHF [9]. Asking human evaluators to provide preferences instead of quantitative scores for each video has the advantage of mitigating the introduction of individual bias [9]. Details of the human feedback collection and Elo scoring are documented in Appendix B.1.
@@ 3.4 Termination
The evolutionary process repeats until it reaches a predetermined number of generations $N$, or when an individual attains a fitness score equivalent to human-level driving performance, as detailed in § 5.1. The final output is the fittest policy $\pi^{\ast}$ corresponding to the highest $\sigma$ in the database $D$. Note that $\pi^{\ast}$ is the best policy in the reward database according to human-assigned fitness scores $\sigma$, but it is not necessarily the optimal policy.
A minimal input-output example of the working of REvolve is shown in Appendix E.
@@ 4 Related Work
Reinforcement Learning and Human Feedback. Due to its complex and hard-to-specify goals [24], autonomous driving presents a challenging problem for reinforcement learning. Autonomous vehicles must maneuver through complex environments and interact with other road users, all the while strictly adhering to traffic regulations. Common techniques to overcome these challenges include learning from demonstrations (LfD) [46] or learning from intervention (LfI) [45]. However, they do not always ensure that the learned behaviors align with human values such as safety, robustness, and efficiency [15]. An alternative is integrating human feedback by using humans as the reward function in the learning process directly [23, 54]. However, this is rather resource-intensive as each rollout needs to be evaluated by a human. REvolve directly integrates human feedback (on the trained policy) into the reward design process. Instead of using humans as reward functions directly, human values, preferences, and knowledge can also be incorporated via so-called human-in-the-loop (HITL) RL. Several techniques exist, for instance, human assessment [9, 23, 54], human demonstration [46], and human intervention [45, 58]. They are particularly beneficial in scenarios like AD, where the reward function is challenging, and the nature of the desired outcome is complex. More recently, reinforcement learning from human feedback (RLHF) [9, 40] has been used successfully in training models like ChatGPT [36], providing guidance and alignment from user preferences. A further comparison of AD approaches with and human feedback is presented in Figure 3, which illustrates how REvolve uses human feedback to align learned policies with human driving standards without requirements for extra reward models, such as in the case of RLHF.
Figure 3: Comparison of REvolve with existing frameworks used to train AD agents. (a.) Expert Designed reward function $R$ to train the AD agent; (b.) learning from demonstrations (LfD) and learning from interventions (LfI), where the agent is trained to imitate the human; (c.) humans acting as reward functions within the training loop, by assessing policy rollouts and providing scalar rewards, which requires significant manual effort; (d.) Reinforcement learning from human feedback (RLHF), where human preference data is used to train an additional (black-box) reward model. This reward model is used as a proxy for human rewards to train the AD policy; (e) Proposed REvolve, which uses GPT-4 as a reward function generator $G$ and evolves them based on (minimal) human feedback on the rollouts sampled from the trained models. This feedback is directly incorporated into the reward design process. REvolve outputs interpretable reward functions, thereby avoiding learning an additional reward model. Here, $\pi\in\Pi$ is a trainable policy in the set of policies $\Pi$. The trainable blocks are denoted by the symbol of the flame.
@@
Reward Design with LLMs. Recent works have explored LLMs for complex search and black-box optimization problems like hyperparameter optimization [28, 62, 65], reward design for RL [25, 30, 61], or robotic control [63]. Kwon et al. [25] use LLMs to directly output binary rewards by inferring user intentions in negotiation games, but this requires numerous queries to LLMs and limits the rewards to sparse binary values. Conversely, Language to Rewards (L2R) [63] and Text2Reward (T2R) [61] utilize LLMs to create dense and interpretable reward functions, improving them with human feedback. L2R, however, restricts its adaptability by employing manually crafted reward templates tailored to specific robot morphologies. In contrast, T2R outputs versatile reward functions in Python code. Eureka [30] further extends T2R by generating multiple reward functions and greedily iterating with the best reward function, and discarding the rest. The greedy search risks premature convergence and resource wastage. Instead, REvolve uses genetic programming to evolve a population of individuals, thus mitigating such risks at no extra computational cost.
@@
Evolutionary Algorithms with LLMs. Implementing evolutionary algorithms with LLMs has been explored in recent works in the context of neural architecture search [34], program search [41], prompt engineering [11, 14], as well as morphology design [27]. In this work, we propose a novel evolutionary framework for reward design in AD with LLMs as intelligent genetic operators.

### v2

# REvolve: Reward Evolution with Large Language Models using Human Feedback

@@ Abstract
Designing effective reward functions is crucial to training reinforcement learning (RL) algorithms. However, this design is non-trivial, even for domain experts, due to the subjective nature of certain tasks that are hard to quantify explicitly. In recent works, large language models (LLMs) have been used for reward generation from natural language task descriptions, leveraging their extensive instruction tuning and commonsense understanding of human behavior. In this work, we hypothesize that LLMs, guided by human feedback, can be used to formulate reward functions that reflect human implicit knowledge. We study this in three challenging settings – autonomous driving, humanoid locomotion, and dexterous manipulation – wherein notions of “good” behavior are tacit and hard to quantify. To this end, we introduce REvolve, a truly evolutionary framework that uses LLMs for reward design in RL. REvolve generates and refines reward functions by utilizing human feedback to guide the evolution process, effectively translating implicit human knowledge into explicit reward functions for training (deep) RL agents. Experimentally, we demonstrate that agents trained on REvolve-designed rewards outperform other state-of-the-art baselines.
@@ 1 Introduction
Recently, LLMs have shown impressive abilities in instruction following (Longpre et al., 2023; Peng et al., 2023), search and black-box optimization (Yang et al., 2024; Zhang et al., 2023; Liu et al., 2024), self-reflection (Madaan et al., 2023; Shinn et al., 2023), and code generation (Chen et al., 2021; Roziere et al., 2023). Owing to their vast training and world knowledge (Wang et al., 2023; Hao et al., 2023; Hazra et al., 2024), they can readily incorporate commonsense priors of human behavior. Together, these abilities make LLMs an ideal candidate for generating rewards from natural language (NL) task descriptions and refining them using language-based feedback loops. Indeed, recent works like Language to Rewards (Yu et al., 2023) and Text2Reward (Xie et al., 2024) have utilized LLMs for this purpose. More significantly, Eureka (Ma et al., 2024a) utilizes LLMs to design dense and interpretable reward functions, achieving improved performance in dexterous manipulation tasks. It does so by generating multiple reward functions and greedily mutating the best candidate over multiple iterations. As such, Eureka resembles more of a greedy iterative search than an evolutionary search, as described in Ma et al. (2024a).
@@
We identify this greedy approach as a significant limitation that risks premature convergence and leads to a waste of resources (by preemptively discarding weaker reward functions). Moreover, Eureka relies on having access to fitness functions to assess the quality of learned behaviors and decide the best reward function. This leads to a chicken-and-egg dilemma – if designing a good fitness measure was feasible, one could arguably design an effective reward function just as easily. This is particularly challenging for complex tasks, such as autonomous driving or dexterous manipulation.
@@
To this end, we introduce Reward Evolve (REvolve), a novel evolutionary framework using LLMs, specifically GPT-4, to output reward functions (as executable Python codes) and evolve them based on human feedback. This allows humans to serve as fitness functions by mapping their preferences and feedback into fitness scores.
In particular, REvolve offers the following novelties:
(1) Using Evolutionary Algorithms (EAs) for Reward Design. Traditional gradient-based optimization is unsuitable for reward design due to the lack of a differentiable cost function. Instead, REvolve utilizes meta-heuristic optimization through EAs (De Jong, 2017) to search for the best candidates. This is done by evolving a population of reward functions using genetic operators like mutation, crossover, and selection – employing LLMs as intelligent operators instead of traditional handcrafted methods. We empirically show that our evolutionary framework considerably outperforms iterative frameworks like Eureka – all without incurring additional computational costs compared to Eureka. Through REvolve, we also demonstrate that EAs, combined with the full range of operations like mutation, crossover, selection, and migration, can be successfully implemented with LLMs.
@@
(2) Utilizing Human Feedback to Guide the Search. Mapping human notions directly into explicit reward functions is challenging; however, humans can still evaluate learned behaviors, effectively acting as fitness functions. Through REvolve, human preferences are directly mapped into fitness scores. Additionally, REvolve incorporates qualitative natural language feedback, offering GPT-4 high-level insights into which aspects to prioritize while refining the reward functions. Together, this ensures that the reward functions – and thus the learned policies – more closely reflect human notions.
@@
(3) Eliminating the Need for Additional Reward Model Training. Unlike reinforcement learning with human feedback (RLHF) (Christiano et al., 2017), which relies on a black-box reward model – trained on massive amounts of curated human preference data – to align LLMs with goals like safety and reliability (OpenAI, 2022), REvolve requires no extra models. REvolve-designed reward functions are interpretable (as executable Python code) compared to RLHF rewards, making it more suitable for safety-critical applications, such as autonomous driving (Christian, 2023). We also include an estimated data requirement comparison in Appendix D.
Figure 1: An overview of REvolve. Given the task (here, autonomous driving) and abstracted environment variables, a reward designer $G$ (LLM) outputs a population of reward functions $R\in\mathcal{R}$, each used to train a policy $\pi(R)$ in simulation. Then, we collect human preferences and natural language (NL) feedback on pairs of policy rollouts $\theta\sim\Theta_{\pi(R)}$ through a user feedback interface. Policy (and thus, corresponding reward function) fitness $\sigma$ is calculated, and the fittest individuals, along with their NL feedback $\lambda$, are refined by $G$. In addition, the process leverages genetic programming for evolution.
@@
Figure 1, provides an overview of REvolve for autonomous driving within a simulated environment.
@@ 2 Problem Formalization
Let us first formally introduce the reward design problem (RDP) (Singh et al., 2009), given by the tuple $\langle M,\mathcal{R},\pi,F\rangle$, where $M=(S,A,T)$ is the world model consisting of state space $S$, action space $A$, and transition function $T$. $\mathcal{R}$ is the set of all reward functions where each $R\in\mathcal{R}$ is a reward function that maps a state-action pair to a real-valued scalar $R:S\times A\rightarrow\mathbb{R}$. The tuple $(M,R)$ constitutes a Markov decision process (Sutton & Barto, 2018). Furthermore, $\pi:S\rightarrow\Delta A$ is a policy that maps a state to a distribution over actions, which is learned using a reward function $R$.
Moreover, the RDP includes a function $F$ that evaluates the fitness (i.e., quality) of the trained policies. Traditionally, $F$ maps policy rollouts, i.e., rollouts of agent-environment interactions, to real-valued scalars. Formally, we denote this by $F:\Omega(\Theta_{\pi})\rightarrow\mathbb{R}$, where $\Theta_{\pi}$ is the random variable modeling the probability distribution of rollouts induced by $\pi$ and $\Omega(\Theta_{\pi})$ is its sample space.
@@
In order to instantiate the RDP tuple for REvolve, we assume that we are given a reward designer $G:\mathcal{L}\rightarrow\mathcal{R}$ that maps from the set of all strings to the set of all reward functions $\mathcal{R}$. The objective now is to find a reward function $R\in\mathcal{R}$ that maximizes the expected fitness score of the policy $\pi$ trained on $R$, which we denote by $\pi(R)$. Formally, we express the optimization problem as follows:
@@
An implicit assumption for reward design is the knowledge of observation and action variables in the environment. For real-world problems, like autonomous driving, the assumption is supported by access to sensor data (such as speed from speedometers and distance covered from odometers) or by employing advanced vision-based models (Kirillov et al., 2023; Cheng et al., 2024; Yuan et al., 2022).
@@ 3 Reward Evolution
As the fitness function $F$ is based on implicit human understanding, it cannot be explicitly formulated and is, therefore, not amenable to gradient-based optimization. Hence, we solve the RDP using a meta-heuristic approach through genetic programming (GP) (Langdon & Poli, 2013). Concretely, REvolve uses GP to output and iteratively refine a population of reward functions (individuals) by applying genetic operations such as mutation, crossover, and migration. In each iteration, we obtain new individuals, which are added to the population based on their fitness. This fitness is assessed through human evaluations of the policies trained with these reward functions – thus guiding the evolutionary search.
Evolutionary algorithms, in general, are considered to be weak methods that perform blind searches and do not exploit domain-specific knowledge (De Jong, 2017). The initialization and perturbations are random, leading to scalability issues (O’Neill et al., 2010). In contrast, LLMs can leverage their commonsense priors to suggest changes that are intelligent and better aligned with the feedback (Gugerty & Olson, 1986). Therefore, we combine the evolutionary search of REvolve with LLMs by using them as reward designers $G$ in the RDP – thus avoiding blindly searching in the solution space.
We follow the island model of evolution (Cantú-Paz et al., 1998), where we maintain multiple ($I$) sub-populations (or islands) in a reward database $D$. Island populations evolve by locally reproducing (mutation and crossover). Periodically, individuals migrate from one island to another. The island model helps enhance genetic diversity across the population and prevent premature convergence in complex search spaces (Cantú-Paz et al., 1998). Following existing works in reward design with LLMs (Xie et al., 2024; Ma et al., 2024a), we represent the reward functions as Python programs that seamlessly integrate with the GP framework. In what follows, we describe the four stages (initialization, reproduction, selection, termination) of our evolutionary framework. We give the formal description in Algorithm 1 – for brevity, we use $\pi$ instead of $\pi(R)$.
@@
10:         mutation $\sim\texttt{Bernoulli}(p_{m})$triangleright Sample operation
@@
25:              $D[P]\leftarrow\texttt{Update}(R,\pi,\sigma,\lambda)$triangleright Add reward to database sub-population
@@ 3.1 Initialization
We start by creating a population of $K$ reward function individuals using GPT-4. This leverages GPT-4’s world knowledge (Hao et al., 2023) and its zero-shot instruction following abilities (Peng et al., 2023). For each individual, we train a policy $\pi$, compute a fitness score $\sigma$, and obtain natural language feedback $\lambda\in\mathcal{L}$. The fitness score calculation and feedback collection are detailed in Section 3.3. Each individual and their respective policy, fitness score, and natural language feedback $(R,\pi,\sigma,\lambda)$ is then randomly assigned to a sub-population and stored in the reward database $D$.
@@ 3.2 Reproduction
Each successive generation of $K$ individuals is created by applying genetic operators (crossover and mutation), on the existing reward database individuals. We use GPT-4 as mutation and crossover operators with appropriate prompts. Due to space constraints, we provide prompt examples for mutation and crossover in Appendix E.
We begin by randomly selecting an operation – mutation with probability $p_{m}$ or crossover with probability $p_{c}=1-p_{m}$. Next, we sample a sub-population $P$ via weighted selection based on the average fitness scores $\sigma^{P}$ of each of the $I$ sub-populations. Subsequently, we sample parents (one for mutation, two for crossover) from $D[P]$ using weighted sampling based on individual fitness scores. These sampled individuals, along with their natural language feedback $\lambda$, are used to prompt GPT-4 to create new individuals. These prompts direct GPT-4 to either modify a single reward component of the parent ($\texttt{Mutate}(G,D[P])$) or combine best performing reward components from two parents ($\texttt{Crossover}(G,D[P])$). Periodically, we enable the migration of individuals between sub-populations to enhance genetic diversity. Reproduction operations are illustrated in Figure 2.
Figure 2: Illustration of how GPT-4 applies mutation and crossover to reward functions. Mutation (left): shows the modification of the ”smoothness reward” component. A red ‘-’ sign indicates the line removed from the parent reward function, while a green ‘+’ sign indicates the line added to the new reward function. Crossover (right): demonstrates how parent reward functions are combined to create a child reward function, incorporating the most effective reward components from each parent.
@@ 3.3 Selection
Newly reproduced individuals (from the reproduction stage) are retained based on fitness scores, following a survival of the fittest approach. A fitness function $F$, tailored to the specific problem, evaluates the quality of each individual. The process involves (1) training a policy $\pi(R)$ until convergence, (2) sampling rollouts $\theta\sim\Theta_{\pi(R)}$, and (3) using the fitness function to obtain a tuple of (real-valued) fitness score and (string) NL feedback $(\sigma,\lambda)\coloneqq F(\theta)$. Individuals are retained if they contribute to an increase in the average fitness of their respective sub-populations $\sigma\geq\sigma^{P}$, where $P$ denotes the sub-population to which the individual is added [^1] To ensure genetic diversity, we add individuals based on surpassing the average and not maximum scores.. This ensures that the average fitness of the sub-population is monotonically increasing. In what follows, we detail the process of obtaining the tuple $(\sigma,\lambda)$ from the fitness function.
To compute the fitness score $\sigma$, we ask human evaluators to judge policy rollouts $\theta$ from different policies in a pairwise fashion. As illustrated in the user feedback interface in Figure 1, these rollouts are presented as video segments $\approx$ 30-40 seconds long. The evaluators choose their preferred segment (ties are also possible). We then apply an Elo rating system (Elo, 1978), which transforms pairwise assessments into fitness scores $\sigma$. This avoids the need to learn a separate reward model, cf. RLHF (Christiano et al., 2017). Asking human evaluators to provide preferences instead of quantitative scores for each video has the advantage of mitigating individual bias (Christiano et al., 2017). Details of the human feedback collection and Elo scoring are documented in Appendix B.1.
@@ 3.4 Termination
The evolutionary process repeats until it reaches a predetermined number of generations $N$, or when an individual attains a fitness score equivalent to human-level driving performance, as detailed in Section 5.3. The final output is the fittest policy $\pi^{\ast}$ corresponding to the highest $\sigma$ in the database $D$. Note that $\pi^{\ast}$ is the best policy in the reward database according to human-assigned fitness scores $\sigma$, but it is not necessarily the optimal policy.
A minimal input-output example of the working of REvolve is shown in Appendix F.
@@ 4 Related Work
Reward Design with LLMs. Recent works have explored LLMs for complex search and black-box optimization problems like hyperparameter optimization (Yang et al., 2024; Zhang et al., 2023; Liu et al., 2024), reward design for RL (Kwon et al., 2023; Xie et al., 2024; Ma et al., 2024a), or robotic control (Yu et al., 2023). Kwon et al. (2023) use LLMs to directly output binary rewards by inferring user intentions in negotiation games, but this requires numerous queries to LLMs and limits the rewards to sparse binary values. Conversely, Language to Rewards (L2R) (Yu et al., 2023) and Text2Reward (T2R) (Xie et al., 2024) utilize LLMs to create dense and interpretable reward functions, improving them with human feedback. L2R, however, restricts its adaptability by employing manually crafted reward templates tailored to specific robot morphologies. In contrast, T2R outputs versatile reward functions in Python code. It generates a single reward function per iteration, checks for improvement from the previous one, and repeats this process, hence performing a greedy search over iterations.
Eureka (Ma et al., 2024a) further extends T2R by generating multiple reward functions and greedily iterating with the best reward function and discarding the rest. This means that by definition (Lehman et al., 2024), Eureka does not qualify as an EA since it lacks crucial features. Namely, maintaining and evolving a population of candidates and employing genetic operators like mutation, crossover, and selection. Eureka only mutates the best candidate and discards the rest. In contrast, REvolve is truly evolutionary, i.e., population-based and using the full range of genetic operators. By preserving genetic diversity, REvolve avoids the pitfalls of premature convergence without incurring additional costs (since both frameworks generate the same number of candidates per generation). Empirically, we show that utilizing the full range of evolutionary operators yields reward functions that lead to better performing agents. Consequently, REvolve significantly outperforms greedy search methods like Eureka and T2R.
@@
Evolutionary Algorithms with LLMs. Implementing EAs with LLMs has been explored in recent works in the context of neural architecture search (Nasir et al., 2024), program search (Romera-Paredes et al., 2024), prompt engineering (Guo et al., 2024; Fernando et al., 2023), as well as morphology design (Lehman et al., 2024). In this work, we propose a novel evolutionary framework for reward design with LLMs as intelligent genetic operators. Compared to the above-mentioned works, which feature limited evolutionary settings, REvolve incorporates a wide range of operations like mutation, crossover, selection, and migration.